OpenStudy (chycora):

6. Choose the equivalent expression. -1.2x + 0.4 + 0.8x + 2x + 0.6 -2x + 3 0.2(-12x + 1) 0.2(8x + 5) 4x + 1.2

OpenStudy (anonymous):

Combine like terms: -1.2x + 0.8x + 2x 0.4 + 0.6 Can you add those?

OpenStudy (anonymous):

then combine the other like terms- -1.2x+2x. can you add those as well?

OpenStudy (anonymous):

Adding terms with the same variable is the same thing as adding regular numbers.. 32b + 9b - 16b 32 + 9 - 16 = 25 So 32b + 9b - 16b = 25b. We just keep the variable.

OpenStudy (anonymous):

So can you add -1.2 + 0.8 + 2? @chycora

OpenStudy (chycora):

yes and hold up real quick

OpenStudy (anonymous):

Okay, tell me what you get.

OpenStudy (chycora):

-1.2x + 0.8x + 2x=4 0.4 + 0.6=1

OpenStudy (anonymous):

Yes, 0.4 + 0.6 = 1 this is correct. but -1.2x + 0.8x + 2x isn't 4..

OpenStudy (chycora):

um 1.6

OpenStudy (anonymous):

Yes, so we have: 1.6x + 1

OpenStudy (anonymous):

So our answer will be C, because when we distribute 0.2 into the parenthesis we get 1.6x + 1.

OpenStudy (chycora):

o ok thanx luv

OpenStudy (chycora):

0.2(8x + 5) so this is the answer right